# GPx4 is bound to peroxidized membranes by a hydrophobic anchor

### v1

@@ 1 Main
Cell death is an essential process during the growth and development of living organisms. The regulation of cell death is crucial for metabolism, immunoregulation and homeostasis. Ferroptosis, coined in 2012, is an iron-dependent and regulated cell death modality induced by phospholipid peroxidation [1]. It is distinctive from other cell death modalities, and many diseases such as diffuse large B cell lymphoma (DLBCL), Huntington’s disease, and Periventricular Leukomalacia (PVL), are related to its misregulation [2, 3]. Using inducers targeting ferroptosis to treat cancers has been in full swing recently because plenty of iron is demanded in the rapid growth of cancer cells, making phospholipids under great peroxidative pressure [4].
@@
Several molecular pathways for ferroptosis regulation have been discovered in recent years [5], among which glutathione peroxidase 4 (GPx4) plays a crucial role in the most classic pathway since it reduces phospholipid peroxides to alcohol by oxidizing glutathiones (GSHs) to glutathione disulfides (GSSGs) [6] (Fig. 1A). Compared with other glutathione peroxidases (GPx) containing selenocysteine (U), GPx4 is the only known one that can treat phospholipid peroxides as redox reaction substrates [7]. Therefore, GPx4 is an essential enzyme that can inhibit ferroptosis by both preventing and repairing cell membranes from the accumulation of peroxides.
@@
To understand the ferroptosis process, computational studies were conducted to study the destructive effect of phospholipid peroxides on cell membranes [8, 9, 10]. Docking calculations showed the likely interaction mode between the GPx4 and lipid bilayers [11], which was in line with the X-ray structure-based analysis, where three key residues, U46, Q81, and W136, form a catalytic triad that should be located at the protein surface toward membranes [12]. Recently, electrostatics was found to be crucial for driving GPx4 interactions with membranes [13] and the R152H mutant could disturb the interaction of the enzyme with membranes [14, 15]. However, a high-resolution picture of GPx4 interacting with membranes is still needed to fully reveal the dynamic anchoring process of GPx4 on peroxidized membranes, which would help develop new regulation strategies.
@@
In this work, by using multiscale molecular dynamics (MD) simulations (Table S1) and experimental assays, we revealed the detailed physical interactions between the enzymatically active U46C mutant [16] of GPx4 and membranes composed of POPC, POPE, POPS, and their peroxidized derivatives (termed HPPC, HPPE and HPPS) (Fig. 1B). Coarse-grained (CG) MD simulations (Fig. S1A–B) were used to study the dynamic binding process of GPx4 on membranes composed of different lipid types (Fig. 1C), which showed a stable and unique binding configuration on PE/PS-rich membranes. Based on the stable binding configuration, all-atom (AA) MD simulations (Fig. S1C–D) were conducted to further evaluate the stability of the binding interface, which also validated L130 was the most important membrane-inserted residue on the GPx4 surface. L130 appeared to serve as a hydrophobic anchor that penetrated deepest into the hydrophobic core of the membrane and stabilized the reaction center of GPx4 toward the membrane surface. Enzyme activity assays and in vitro cell experiments provided further validations, highlighting the important role of the hydrophobic residue L130 in stabilizing the anchoring configuration of GPx4 on membranes, which was also proved by in silico simulations.
@@
The difference in the binding affinity may arise from the different types of the head groups of these phospholipids, representing their different features in physical chemistry. Although the head groups of PC/PE/PS are all represented by one bead in the Martini force field, they have different non-bonded interaction parameters, with POPC being less attractive or more repulsive than POPE and POPS [18]. This may correspond to the fact that the PC head group is larger than PE or PS in the atomic model, so that the head group of POPC can lead to a more compact hydrophilic wall and less space for GPx4 to be inserted and reach the hydrophobic core of the lipid bilayer. Notably, the binding of GPx4 to the membrane with peroxidized phospholipids showed high affinity as well, where the GPx4 was never detached from the membrane surface once it was stably bound in our simulation (Fig. 2D).
@@
As GPx4 can inhibit ferroptosis by binding and repairing peroxidized cell membranes, we utilized peroxidized liposomes to assess the activity of GPx4 and to evaluate whether mutations at the hydrophobic L130 residue diminish GPx4’s activity to repair membranes. Liposomes were prepared with a phospholipid mixture and enclosed 5-Carboxyfluorescein [20], followed by incubation with purified ALOX15, which was sufficient to cause leakage from the liposomes. The fluorescein within the liposomes exhibits a weak signal and the signal intensity will increase significantly if it is released from the liposomes. Here, we also used the enzymatically active GPx4-U46C mutant [16] to substitute for the wild type of GPx4. GPx4 addition suppressed ALOX15-induced liposome leakage, confirming its membrane repair activity (Fig. S8). Substitution of the hydrophobic L130 residue with polar residues (e.g., L130S and L130Q) markedly reduced GPx4’s membrane repair capability (Fig. 5A, left panel). Importantly, these mutations did not decrease the activity of GPx4 in simple soluble substrates in solution (Fig. 5A, right panel). Given that the expression levels of GPx4 and its mutants were comparable (Fig. S9A), these results indicated that the weakening of membrane repair ability stems from impaired membrane binding of the mutated GPx4.
@@
2 Materials and Methods
2.1 System setup
@@
The structure of GPx4 was downloaded from the Protein Data Bank (PDB ID: 2OBI), and then the CG model was generated by martinize.py [21]. It should be noted that the structure of GPx4-U46C instead of GPx4 was used in our MD simulations due to the lack of parameters for Se in the available force fields, but GPx4-U46C is still enzymatically active [16] and the results were not expected to be significantly affected for our purpose because this would not alter the likely interaction interface between the GPx4 and membranes. In the virtual mutation validation, GPx4 U46C/L130A, U46C/L130S and U46C/L130Q were constructed with Chimera [22], and the rotamer library Dunbrack 2010 was used for modeling [23].
@@
The model membranes used in our simulations were composed of 2-oleoyl-1-palmitoyl-sn-glycero-3-phosphocholine (POPC), 1-palmitoyl-2-oleoyl-sn-glycero-3-phosphoethanolamine (POPE), 1-palmitoyl-2-oleoyl-sn-glycero-3-phospho-L-serine (POPS) and their corresponding peroxides, named HPPC, HPPE and HPPS respectively (Fig. 1B). Parameter files of POPC, POPE, and POPS were readily available, while HPPC, HPPE, and HPPS were new types of phospholipids that needed to be parameterized. Fabrice Thalmann et al. developed the CG model and parameter file of HPPC (i.e. HP-POPC) [24] within the Martini2.2 force field [18], and the parameters for HPPE and HPPS were created in the same way. The AA parameter files of HPPC, HPPE, and HPPS were first generated by the module Ligand Reader & Modeler [25] in CHARMM-GUI [26], and then the atom types within these files were further checked and adjusted to ensure that the parameters of HPPC, HPPE, and HPPS were consistent with those of POPC, POPE, and POPS, except for the additional peroxidized group and the connected carbon atom. Four types of phospholipid bilayers were constructed for our simulations: a pure POPC bilayer, a mixed bilayer consisting of POPC and POPE (POPC: POPE = 1:1), a mixed bilayer consisting of POPC, POPE, and POPS (POPC: POPE: POPS = 1:2:2), and a mixed POPC/POPE/POPS bilayer containing peroxidized phospholipids (POPC: POPE: POPS: HPPC: HPPE: HPPS = 1:2:2:1:2:2). All the simulated bilayers are symmetric, that is, for the same membrane, the phospholipid compositions of the inner and outer leaflets are the same. The phospholipid composition ratio of the POPC/POPE/POPS membrane is set close to the inner bilayer of plasma membranes [27, 28, 29, 30, 17, 31]. The CG simulation systems were generated with insane.py [32] and solved in a water box of 12 nm × 12 nm × 15 nm (Fig. S1A). The script backward.py [33] was used to build AA models (Fig. S1C) from the last frame of the CG simulations.
2.2 Molecular dynamics simulations
@@
CG simulations were performed with the GROMACS 2021.2 package [34] using the Martini2.2 force field with an elastic network [18]. To achieve energy minimization, the steepest descent algorithm was used for thousands of steps and the resulting maximum force on each atom was less than 10 kJ mol-1 nm-2. Then, the systems underwent a 500-ps equilibration to reach 310 K with position restraints on the backbone beads (force constant 1000 kJ mol-1 nm-2) in the NVT ensemble. Then the production simulations were performed under the NPT ensemble and the temperature and the pressure were maintained at 310 K and 1.0 bar by using the V-rescale algorithm [35] with a time constant of 1.0 ps and the Berendsen algorithm [36] with a time constant of 10.0 ps, respectively. The Berendsen barostat was used to adapt to the possible changes in box size caused by the insertion of GPx4 into the membrane. In all the simulations, the periodic boundary conditions were applied and Newton’s equations of motion were integrated at intervals of 10 fs. Electrostatics were calculated with a cutoff of 1.1 nm and a relative dielectric constant of 15, while the permittivity was set to infinity beyond the cutoff.
@@
AA simulations were performed with the GROMACS 2021.2 package [34] and conducted with the CHARMM36m force field [26]. 0.15 mol/L NaCl was added to the AA system. Before AA production simulations, an energy minimization with thousands of steps, a 500-ps NVT equilibration, and a 50-ns semi-isotropic NPT equilibration were conducted to reach a temperature of 310 K and a pressure of 1.0 bar, and position restraints were applied on C atoms of GPx4, with a force constant of 1000 kJ mol-1 nm-2. The production simulations were under the semi-isotropic NPT ensemble without any restraints, in which the V-rescale algorithm [37] with a time constant of 0.5 ps and the Parrinello-Rahman algorithm [38] with a time constant of 1.0 ps were used to maintain the temperature at 310 K and the pressure at 1.0 bar, respectively. In all of AA simulations, the periodic boundary conditions were used and the time step was set to 2 fs. The short-range cut-offs were 1.2 nm for electrostatic and van der Waals (vdW) interactions, with a switch function turning off the vdW interactions from 1.0 to 1.2 nm. The long-range electrostatic interactions were calculated with the particle mesh Ewald (PME) method [39].
2.3 Experiments
@@
ALOX15, GPx4 and GPx4 mutants were expressed using Escherichia coli expression system as previously described[42, 43]. In brief, ALOX15 with a His-tag was purified using a nickel-nitrilotriacetic acid column (HisTrap HP, GE Healthcare) and an anion-exchange column (Q Trap HP, GE Healthcare). GPx4 and mutants were fused with His-tag and purified using a nickel-nitrilotriacetic acid column at 4 °C. The eluted enzyme was further subjected to dialysis for 12 h [100 mM Tris-HCl (pH 7.4), 1 mM DTT, and 20% (v/v) glycerol] to remove the bulk of the imidazole. The final purity of the enzyme was >95% as confirmed by sodium dodecyl sulfate–polyacrylamide gel electrophoresis. Protein concentrations were measured using Nanodrop 2000 (Thermo Scientific).
@@
Liposomes were prepared as previously described [44]. 16:0 PC (DPPC), and 18:0-20:4 PE (SAPE) were obtained from Cayman Chemical. The phospholipids (80% DPPC and 20% SAPE) were dissolved in chloroform as a 10 mg/ml stock solution. The phospholipid films were obtained by mixing a 100 L stock solution and 500 L of chloroform in round-bottomed 10 mL flasks and evaporating the solvent using a vacuum rotary evaporator. The lipid films were hydrated with 500 L of buffer L (50 mM PBS, pH 7.3, 100 mM NaCl) and the liposomes were subjected to 20 rounds of extrusion through a 100 nm polycarbonate membrane using a mini-extruder (Avanti). To prepare the 5-Carboxyfluorescein encapsulated liposomes, the lipid films were hydrated with 500 L of buffer CL (50 mM PBS, pH 7.3, 100 mM NaCl, 20 mM 5-Carboxyfluorescein). After performing the extrusion, the liposomes were purified by a desalting column to remove any external 5-Carboxyfluorescein, after which they were re-suspended in 500 L of buffer L. All the liposomes were stored at 4 °C and used within 24 h.
@@
The reactivity of ALOX15 with phospholipids in liposomes has been reported [45]. For the liposome leakage assay, 10 L of the liposomes were diluted into 100 L of buffer LA [100 mM Tris-HCl (pH 7.4), 3 mM glutathione, 100 mM NaCl]. When 50 M of ALOX15 and 50 M of GPx4 were added to the reaction mixture, the fluorescence signal was measured by a microplate reader. The fluorescence signal was then recorded at 10-s intervals and the relative activity of different mutants was compared by calculating the change of fluorescence signal intensity within ten minutes.

### v2

@@ Abstract
Cell death is an essential process during the growth and development of living organisms. The regulation of cell death is crucial for metabolism, immunoregulation and homeostasis. Ferroptosis, coined in 2012, is an iron-dependent and regulated cell death modality induced by phospholipid peroxidation 1. It is distinctive from other cell death modalities, and many diseases such as diffuse large B cell lymphoma (DLBCL), Huntington’s disease, and Periventricular Leukomalacia (PVL), are related to its misregulation 2, 3. Using inducers targeting ferroptosis to treat cancers has been in full swing recently because plenty of iron is demanded in the rapid growth of cancer cells, making phospholipids under great peroxidative pressure 4.
@@
Several molecular pathways for ferroptosis regulation have been discovered in recent years 5, among which glutathione peroxidase 4 (GPx4) plays a crucial role in the most classic pathway since it reduces phospholipid peroxides to alcohol by oxidizing glutathiones (GSHs) to glutathione disulfides (GSSGs) 6 (Fig. 1A). Compared with other glutathione peroxidases (GPx) containing selenocysteine (U), GPx4 is the only known one that can treat phospholipid peroxides as redox reaction substrates 7. Therefore, GPx4 is an essential enzyme that can inhibit ferroptosis by both preventing and repairing cell membranes from the accumulation of peroxides.
@@
To understand the ferroptosis process, computational studies were conducted to study the destructive effect of phospholipid peroxides on cell membranes 8, 9, 10. Docking calculations showed the likely interaction mode between the GPx4 and lipid bilayers 11, which was in line with the X-ray structure-based analysis, where three key residues, U46, Q81, and W136, form a catalytic triad that should be located at the protein surface toward membranes 12. Recently, electrostatics was found to be crucial for driving GPx4 interactions with membranes 13 and the R152H mutant could disturb the interaction of the enzyme with membranes 14, 15. However, a high-resolution picture of GPx4 interacting with membranes is still needed to fully reveal the dynamic anchoring process of GPx4 on peroxidized membranes, which would help develop new regulation strategies.
@@
In this work, by using multiscale molecular dynamics (MD) simulations (Table S1) and experimental assays, we revealed the detailed physical interactions between the enzymatically active U46C mutant 16 of GPx4 and membranes composed of POPC, POPE, POPS, and their peroxidized derivatives (termed HPPC, HPPE and HPPS) (Fig. 1B). Coarse-grained (CG) MD simulations (Fig. S1A–B) were used to study the dynamic binding process of GPx4 on membranes composed of different lipid types (Fig. 1C), which showed a stable and unique binding configuration on PE/PS-rich membranes. Based on the stable binding configuration, all-atom (AA) MD simulations (Fig. S1C–D) were conducted to further evaluate the stability of the binding interface, which also validated L130 was the most important membrane-inserted residue on the GPx4 surface. L130 appeared to serve as a hydrophobic anchor that penetrated deepest into the hydrophobic core of the membrane and stabilized the reaction center of GPx4 toward the membrane surface. Enzyme activity assays and in vitro cell experiments provided further validations, highlighting the important role of the hydrophobic residue L130 in stabilizing the anchoring configuration of GPx4 on membranes, which was also proved by in silico simulations.
@@
The difference in the binding affinity may arise from the different types of the head groups of these phospholipids, representing their different features in physical chemistry. Although the head groups of PC/PE/PS are all represented by one bead in the Martini force field, they have different non-bonded interaction parameters, with POPC being less attractive or more repulsive than POPE and POPS 18. This may correspond to the fact that the PC head group is larger than PE or PS in the atomic model, so that the head group of POPC can lead to a more compact hydrophilic wall and less space for GPx4 to be inserted and reach the hydrophobic core of the lipid bilayer. Notably, the binding of GPx4 to the membrane with peroxidized phospholipids showed high affinity as well, where the GPx4 was never detached from the membrane surface once it was stably bound in our simulation (Fig. 2D).
@@
As GPx4 can inhibit ferroptosis by binding and repairing peroxidized cell membranes, we utilized peroxidized liposomes to assess the activity of GPx4 and to evaluate whether mutations at the hydrophobic L130 residue diminish GPx4’s activity to repair membranes. Liposomes were prepared with a phospholipid mixture and enclosed 5-Carboxyfluorescein 20, followed by incubation with purified ALOX15, which was sufficient to cause leakage from the liposomes. The fluorescein within the liposomes exhibits a weak signal and the signal intensity will increase significantly if it is released from the liposomes. Here, we also used the enzymatically active GPx4-U46C mutant 16 to substitute for the wild type of GPx4. GPx4 addition suppressed ALOX15-induced liposome leakage, confirming its membrane repair activity (Fig. S8). Substitution of the hydrophobic L130 residue with polar residues (e.g., L130S and L130Q) markedly reduced GPx4’s membrane repair capability (Fig. 5A, left panel). Importantly, these mutations did not decrease the activity of GPx4 in simple soluble substrates in solution (Fig. 5A, right panel). Given that the expression levels of GPx4 and its mutants were comparable (Fig. S9A), these results indicated that the weakening of membrane repair ability stems from impaired membrane binding of the mutated GPx4.
@@
Materials and Methods
System setup
@@
The structure of GPx4 was downloaded from the Protein Data Bank (PDB ID: 2OBI), and then the CG model was generated by martinize.py 21. It should be noted that the structure of GPx4-U46C instead of GPx4 was used in our MD simulations due to the lack of parameters for Se in the available force fields, but GPx4-U46C is still enzymatically active 16 and the results were not expected to be significantly affected for our purpose because this would not alter the likely interaction interface between the GPx4 and membranes. In the virtual mutation validation, GPx4 U46C/L130A, U46C/L130S and U46C/L130Q were constructed with Chimera 22, and the rotamer library Dunbrack 2010 was used for modeling 23.
@@
The model membranes used in our simulations were composed of 2-oleoyl-1-palmitoyl-sn-glycero-3-phosphocholine (POPC), 1-palmitoyl-2-oleoyl-sn-glycero-3-phosphoethanolamine (POPE), 1-palmitoyl-2-oleoyl-sn-glycero-3-phospho-L-serine (POPS) and their corresponding peroxides, named HPPC, HPPE and HPPS respectively (Fig. 1B). Parameter files of POPC, POPE, and POPS were readily available, while HPPC, HPPE, and HPPS were new types of phospholipids that needed to be parameterized. Fabrice Thalmann et al. developed the CG model and parameter file of HPPC (i.e. HP-POPC) 24 within the Martini2.2 force field 18, and the parameters for HPPE and HPPS were created in the same way. The AA parameter files of HPPC, HPPE, and HPPS were first generated by the module Ligand Reader & Modeler 25 in CHARMM-GUI 26, and then the atom types within these files were further checked and adjusted to ensure that the parameters of HPPC, HPPE, and HPPS were consistent with those of POPC, POPE, and POPS, except for the additional peroxidized group and the connected carbon atom. Four types of phospholipid bilayers were constructed for our simulations: a pure POPC bilayer, a mixed bilayer consisting of POPC and POPE (POPC: POPE = 1:1), a mixed bilayer consisting of POPC, POPE, and POPS (POPC: POPE: POPS = 1:2:2), and a mixed POPC/POPE/POPS bilayer containing peroxidized phospholipids (POPC: POPE: POPS: HPPC: HPPE: HPPS = 1:2:2:1:2:2). All the simulated bilayers are symmetric, that is, for the same membrane, the phospholipid compositions of the inner and outer leaflets are the same. The phospholipid composition ratio of the POPC/POPE/POPS membrane is set close to the inner bilayer of plasma membranes 27, 28, 29, 30, 17, 31. The CG simulation systems were generated with insane.py 32 and solved in a water box of 12 nm × 12 nm × 15 nm (Fig. S1A). The script backward.py 33 was used to build AA models (Fig. S1C) from the last frame of the CG simulations.
Molecular dynamics simulations
@@
CG simulations were performed with the GROMACS 2021.2 package 34 using the Martini2.2 force field with an elastic network 18. To achieve energy minimization, the steepest descent algorithm was used for thousands of steps and the resulting maximum force on each atom was less than 10 kJ mol-1 nm-2. Then, the systems underwent a 500-ps equilibration to reach 310 K with position restraints on the backbone beads (force constant 1000 kJ mol-1 nm-2) in the NVT ensemble. Then the production simulations were performed under the NPT ensemble and the temperature and the pressure were maintained at 310 K and 1.0 bar by using the V-rescale algorithm 35 with a time constant of 1.0 ps and the Berendsen algorithm 36 with a time constant of 10.0 ps, respectively. The Berendsen barostat was used to adapt to the possible changes in box size caused by the insertion of GPx4 into the membrane. In all the simulations, the periodic boundary conditions were applied and Newton’s equations of motion were integrated at intervals of 10 fs. Electrostatics were calculated with a cutoff of 1.1 nm and a relative dielectric constant of 15, while the permittivity was set to infinity beyond the cutoff.
@@
AA simulations were performed with the GROMACS 2021.2 package 34 and conducted with the CHARMM36m force field 26. 0.15 mol/L NaCl was added to the AA system. Before AA production simulations, an energy minimization with thousands of steps, a 500-ps NVT equilibration, and a 50-ns semi-isotropic NPT equilibration were conducted to reach a temperature of 310 K and a pressure of 1.0 bar, and position restraints were applied on C atoms of GPx4, with a force constant of 1000 kJ mol-1 nm-2. The production simulations were under the semi-isotropic NPT ensemble without any restraints, in which the V-rescale algorithm 37 with a time constant of 0.5 ps and the Parrinello-Rahman algorithm 38 with a time constant of 1.0 ps were used to maintain the temperature at 310 K and the pressure at 1.0 bar, respectively. In all of AA simulations, the periodic boundary conditions were used and the time step was set to 2 fs. The short-range cut-offs were 1.2 nm for electrostatic and van der Waals (vdW) interactions, with a switch function turning off the vdW interactions from 1.0 to 1.2 nm. The long-range electrostatic interactions were calculated with the particle mesh Ewald (PME) method 39.
Experiments
@@
ALOX15, GPx4 and GPx4 mutants were expressed using Escherichia coli expression system as previously described42, 43. In brief, ALOX15 with a His-tag was purified using a nickel-nitrilotriacetic acid column (HisTrap HP, GE Healthcare) and an anion-exchange column (Q Trap HP, GE Healthcare). GPx4 and mutants were fused with His-tag and purified using a nickel-nitrilotriacetic acid column at 4 °C. The eluted enzyme was further subjected to dialysis for 12 h [100 mM Tris-HCl (pH 7.4), 1 mM DTT, and 20% (v/v) glycerol] to remove the bulk of the imidazole. The final purity of the enzyme was >95% as confirmed by sodium dodecyl sulfate–polyacrylamide gel electrophoresis. Protein concentrations were measured using Nanodrop 2000 (Thermo Scientific).
@@
Liposomes were prepared as previously described 44. 16:0 PC (DPPC), and 18:0-20:4 PE (SAPE) were obtained from Cayman Chemical. The phospholipids (80% DPPC and 20% SAPE) were dissolved in chloroform as a 10 mg/ml stock solution. The phospholipid films were obtained by mixing a 100 L stock solution and 500 L of chloroform in round-bottomed 10 mL flasks and evaporating the solvent using a vacuum rotary evaporator. The lipid films were hydrated with 500 L of buffer L (50 mM PBS, pH 7.3, 100 mM NaCl) and the liposomes were subjected to 20 rounds of extrusion through a 100 nm polycarbonate membrane using a mini-extruder (Avanti). To prepare the 5-Carboxyfluorescein encapsulated liposomes, the lipid films were hydrated with 500 L of buffer CL (50 mM PBS, pH 7.3, 100 mM NaCl, 20 mM 5-Carboxyfluorescein). After performing the extrusion, the liposomes were purified by a desalting column to remove any external 5-Carboxyfluorescein, after which they were re-suspended in 500 L of buffer L. All the liposomes were stored at 4 °C and used within 24 h.
@@
The reactivity of ALOX15 with phospholipids in liposomes has been reported 45. For the liposome leakage assay, 10 L of the liposomes were diluted into 100 L of buffer LA [100 mM Tris-HCl (pH 7.4), 3 mM glutathione, 100 mM NaCl]. When 50 M of ALOX15 and 50 M of GPx4 were added to the reaction mixture, the fluorescence signal was measured by a microplate reader. The fluorescence signal was then recorded at 10-s intervals and the relative activity of different mutants was compared by calculating the change of fluorescence signal intensity within ten minutes.
@@ Acknowledgements
| System ID | Protein | Membrane | Bead numbers | Box size (nm3) | Simulation time (μs) | Replicate numbers |
| CG-1 | GPx4-U46C | POPC 100% | 18264 | 50 | 1 | |
| CG-2 | GPx4-U46C | POPC 50% POPE 50% | 18292 | 50 | 1 | |
| CG-3 | GPx4-U46C | POPC 20% POPE 40% POPS 40% | 18175 | 50 | 1 | |
| CG-4 | GPx4-U46C | POPC 100% | 18175 | 10 | 3 (mutate PE and PS to PC based on 50-μs result of CG-3) | |
| CG-5 | GPx4-U46C | peroxidized PC/PE/PS membrane | 18693 | 50 | 1 | |
| CG-6 | GPx4-U46C | peroxidized PC/PE/PS membrane | 18693 | 10 | 66 (six different initial configurations) | |
| AA-1 | GPx4-U46C | peroxidized PC/PE/PS membrane | 211625 | 0.5 | 3 (L130 inserted in the membrane) | |
| AA-2 | GPx4-U46C | peroxidized PC/PE/PS membrane | 211621 | 0.5 | 3 (L130 was close to the membrane surface) | |
| AA-3 | GPx4-U46C/L130S | peroxidized PC/PE/PS membrane | 222617 | 0.5 | 3 (S130 inserted in the membrane) | |
| CG-7 | GPx4-U46C | peroxidized PC/PE/PS membrane | 18693 | 10 | 3 | |
| CG-8 | GPx4-U46C/L130A | peroxidized PC/PE/PS membrane | 18692 | 10 | 3 | |
| CG-9 | GPx4-U46C/L130S | peroxidized PC/PE/PS membrane | 18693 | 10 | 3 | |
| CG-10 | GPx4-U46C/L130Q | peroxidized PC/PE/PS membrane | 18693 | 10 | 3 |